# First principle prediction of structural distortions in the cuprates and their impact on the electronic structure

### v2

@@ Abstract
Materials-realistic microscopic theoretical descriptions of copper-based superconductors are challenging due to their complex crystal structures combined with strong electron interactions. Here, we demonstrate how density functional theory can accurately describe key structural, electronic, and magnetic properties of the normal state of the prototypical cuprate Bi2Sr2CaCu2O8+x (Bi-2212). We emphasize the importance of accounting for energy-lowering structural distortions, which then allows us to: (a) accurately describe the insulating antiferromagnetic (AFM) ground state of the undoped parent compound (in contrast to the metallic state predicted by previous ab initio studies); (b) identify numerous low-energy competing spin and charge stripe orders in the hole-overdoped material nearly degenerate in energy with the AFM ordered state, indicating strong spin fluctuations; (c) predict the lowest-energy hole-doped crystal structure including its long-range structural distortions and oxygen dopant positions that match high-resolution scanning microscopy measurements; and (d) describe electronic bands near the Fermi energy with flat antinodal dispersions and Fermi surfaces that in agreement with angle-resolved photoemission spectroscopy (ARPES) measurements and provide a clear explanation for the structural origins of the so-called “shadow bands”. We also show how one must go beyond band theory and include fully dynamic spin fluctuations via a many-body approach when aiming to make quantitative predictions to measure the ARPES spectra in the overdoped material.
@@ I Introduction
The cuprate superconductors continue to be a fascinating and actively researched class of materials. In addition to their superconducting phases, their normal state has attracted broad research interest due to a wide range of unusual properties. Understanding the physical origin of the AFM insulating phase Kastner et al. (1998); Ho et al. (2001), the pseudogap Warren et al. (1989); Homes et al. (1993); Loeser et al. (1996); Ding et al. (1996), the flat antinodal dispersion Andersen et al. (1994); Liechtenstein et al. (1996); He et al. (2021); Li et al. (2021), the strange metallicity Gurvitch and Fiory (1987); Martin et al. (1990); Cooper et al. (2009); Greene et al. (2020); Ayres et al. (2021), and the presence of quantum critical fluctuations Dai et al. (2001); Le Tacon et al. (2011); Dean et al. (2013); Keimer et al. (2015) in the normal state can provide important insights into the underlying mechanisms that can give rise to the superconductivity. Despite plenty of proposed possible mechanisms, including competing orders Kivelson et al. (1998); Lee et al. (2006); Chang et al. (2012) and preformed pairs Emery and Kivelson (1995); Corson et al. (1999); Li et al. (2010); Dubroka et al. (2011), the physical origin of this complicated normal state is still unclear.
@@
A comprehensive description of the normal state is challenging due to the strong electronic interactions combined with the structural complexity of typical doped cuprates. While the effect of strong electronic interactions has been extensively studied using accurate many-body methods such as density matrix renormalization group (DMRG) White (1992); Schollwöck (2005) or quantum Monte Carlo (QMC) Ceperley and Alder (1986); Foulkes et al. (2001), these studies are often based on idealized effective model Hamiltonian, where the hopping strengths are averaged and symmetrized for simplicity Cui et al. (2020); White and Scalapino (2015); Huang et al. (2017). However, realistic structural distortions in cuprates can result in significant changes to the materials such as additional symmetry breaking Gao et al. (1988); Mans et al. (2006). In particular, the structural distortion in cuprates can greatly modify the superconducting gap Andersen et al. (2007); McElroy et al. (2005) and local pairing interactions Foyevtsova et al. (2009); Gastiasoro and Andersen (2018); Nunner et al. (2005); Maśka et al. (2007); Khaliullin et al. (2010); Litak (2009); Rømer et al. (2013); Johnston et al. (2009). These properties are missing in the symmetrized effective Hamiltonians. The inclusions of complex lattice distortions of the native material or as introduced by dopants and impurities necessitate a realistic and detailed understanding of the materials from first principles.
@@
Density functional theory Hohenberg and Kohn (1964); Kohn and Sham (1965) (DFT) offers a potent foundational method for investigating the ground-state properties of materials from first principles. For cuprates, DFT has played a pivotal role in constructing effective model Hamiltonians Hybertsen et al. (1989); Martin (1996); Markiewicz et al. (2005); Korshunov et al. (2005); Delannoy et al. (2009); Hanke et al. (2010); Hirayama et al. (2019); Ohgoe et al. (2020); Nilsson et al. (2019); Cui et al. (2022, 2023); Morée et al. (2022); Schmid et al. (2023). These works studied the high-symmetry cuprate crystal structure using DFT and extracted the low-energy effective model for further quantum many-body studies. However, to ensure the accuracy and realism of these models, it is essential that the DFT calculations capture correctly both the structural properties and the predominant electronic properties of the ground state. Recent DFT studies on transition metal oxides have highlighted the significance of allowing energy-lowering structural distortions to achieve high-quality predictions of materials properties, such as band gaps at eV scale Varignon et al. (2019); Wang et al. (2021a); Zunger (2022).
@@
Bismuth strontium calcium copper oxide Bi2Sr2CaCu2O8+x (BSCCO or Bi-2212) Maeda et al. (1988) is one of the most intensively studied cuprates and is the focus of in this work. However, prior DFT studies faced challenges due to its intricate structural distortions and superlattice modulations Petricek et al. (1990); Levin et al. (1994); Slezak et al. (2008); He et al. (2008). We demonstrate that by employing modern DFT exchange-correlation functionals and providing an accurate description of the crystalline structure including energy-lowering lattice distortions, we can directly describe the antiferromagnetic insulating ground state of undoped Bi-2212 ($x=0$), the experimentally-observed crystal structure of hole-doped Bi-2212 ($x\approx 0.25$), the presence of competing magnetic and charge stripe orders, as well as crucial details of photoemission spectra such as “shadow bands”. This means that based on DFT calculations, we can correctly ascribe certain experimental observations to specific structural motifs (e.g., shadow bands) while simultaneously helping build microscopically well-justified model Hamiltonians that can be used to compute the effects of strong electron correlations.
@@ II Undoped system
The undoped Bi-2212 possesses a bilayer crystal structure as depicted in Fig. 1(a)(b), where each bilayer comprises two CuO layers separated by one Ca layer and sandwiched between SrO and BiO layers. The smallest unit cell contains 30 atoms and crystallizes in the tetragonal I4/mmm space group, with only one Cu atom in each CuO layer Liang et al. (1988); Jain et al. (2013); mat . However, this small unit cell leads to a false non-magnetic ground state due to the artificial assumption of translational invariance of the Cu local moments Yelpo et al. (2021). In addition, the tetragonal space group of this small unit cell is inconsistent with the orthorhombicity observed in experiments Zeljkovic et al. (2012). Here, we start with a larger supercell (60 atoms/cell) with two Cu atoms in each CuO layer, which allows for structural orthorhombicity and the spontaneous symmetry breaking of the magnetic local moments. We ignore the superlattice modulation distortion for now and focus on the distortions in BiO layers. As we will show in the following, the distortions in BiO layers are crucial to raising the antibonding Bi-O band to higher energy.
For simplicity, we first compare different distortion patterns in bulk calculations. After finding the most energetically favorable distortions, we then perform slab calculation to compare with available undoped Bi-2212 experiments such as STM and thin film transports. Because the inter-bilayer van der Waals interactions in Bi-2212 are weak, the slab calculations show very small differences with the bulk results. This is also the reason why many prior studies only focused on slab calculations Wang et al. (2005); He et al. (2006); Foyevtsova et al. (2010).
@@
To improve the theoretical description, allowing crystal structural distortions in the BiO layers can further lower the total energy of the calculation. Using the conjugate gradient algorithm for structural relaxation, we optimize the structures by initially lowering the symmetry in the BiO layers manually. Fig. 2(a)(b) depicts two typical stable or meta-stable BiO layer patterns whose ground states are both G-AFM ordered on the CuO2 planes. These two distortions have been studied separately in prior studies Bellini et al. (2004); Lin et al. (2006); Nokelainen et al. (2020). Here we will compare the effects of different distortion patterns on electronic structures and demonstrate how the distortions help raise the BiO antibonding band.
@@
Fig. 2(b) shows an orthorhombic pattern, characterized by the lowest symmetry among all three structures presented in Fig. 1 and 2. This pattern further breaks the mirror symmetry of the crystal, resulting in an orthorhombic lattice consistent with observations in experiments Zeljkovic et al. (2012). This orthorhombic pattern exhibits the lowest AFM ground state energy, 0.5 eV/Cu lower than the high-symmetry case. It also features the largest local Cu magnetic moments of ±$\pm 0.53\mu_{B}$ as well as the largest BiO bonding gap size of about 1.9 eV among all three Bi-O structural motifs mentioned above.
@@
This microscopic picture is consistent with a structural analysis: the Bi-O bond lengths in the high-symmetry, zigzag, and orthorhombic structures are 2.65, 2.28, and 2.17Å, respectively. In general, a shorter bond length implies a larger hopping element between two atoms. We substantiate this picture quantitatively by computing the tight-binding Kohn-Sham Hamiltonian on the maximally localized Wannier basis Marzari and Vanderbilt (1997) extracted from our DFT calculations using Wannier90 Pizzi et al. (2020). The projected Wannier orbitals encompass the $p$-orbitals of Bi and O, as well as the $d$-orbitals of Cu, which are sufficient to describe the bands near the Fermi level Jin and Ismail-Beigi . Our calculations reveal that the Bi-O hopping (tunneling) matrix element for nearest-neighbor in-plane Bi-O pairs is about 1.0, 1.8, and 2.3 eV for the high-symmetry, zigzag, and orthorhombic structures, respectively, which is consistent with the behavior of the gaps. Notably, the hopping strength in the orthorhombic structure is the only one sufficiently large to open a gap and raise the Bi electron pockets (originating from the anti-bonding BiO bands) above the Fermi level, which leads to the insulating ground state.
@@
We conclude this section by comparing our results with experiments and prior theories. Experimentally, bulk Bi-2212 crystals are conducting due to hole hoping from inevitable excess oxygen $(x>0)$, so comparing our undoped calculations to experiments on the bulk is not fruitful. However, undoped thin films of Bi-2212 ($x=0$) have been realized experimentally: scanning tunneling microscopy (STM) Ruan et al. (2016); Kowalski et al. (2021); Wang et al. (2023) and transport Yu et al. (2019); Wang et al. (2021b) experiments have found an AFM insulating ground state in undoped thin films Bi-2212. (About $x\approx 6\%$ hole-doping is required to turn the insulating Bi-2212 film into a superconductor Yu et al. (2019); Wang et al. (2021b).) Theoretically, some aspects of the undoped ground state have not been described well by prior DFT work which may have hampered further theoretical studies of the effects of doping and other perturbations on the system at a microscopic level. For example, prior DFT works were unable to simultaneously reproduce the AFM order and the insulating behavior of undoped Bi-2212 whether modeling the bulk Yelpo et al. (2021); Lin et al. (2006); Nokelainen et al. (2020), thin film or or slabs Wang et al. (2005); He et al. (2006); Foyevtsova et al. (2010), or when using larger more realistic unit cells that include the observed crystal modulation He et al. (2008); Fan and Zeng (2011). The prior work found undoped Bi-2212 to be metallic and “self-doped” due to Bi electron pockets at the Fermi surface (pockets that have not been seen in experiments). Some prior works Lin et al. (2006); He et al. (2006); Foyevtsova et al. (2010) resorted to additional “manual” hole doping to eliminate the finite Bi density of states: while this is a sensible workaround, it would be preferable to have a calculation of the undoped material reflect the insulating experimental system.
@@ III Oxygen-doped system
Figure 4: (a) DFT predicted crystal structure of Bi2Sr2CaCu2O8.25, where the hole doping level is $x=0.25$. Sr, Ca, Cu, Bi, O, and dopant O atoms are marked by green, gray, blue, purple, red, and black balls, respectively. Red arrows further highlight the O dopants. The black solid square in the crystal structure marks the 244-atom unit cell of the doped crystal. The black dashed rectangle highlights the CuO layers in one of the bilayers in the unit cell. (b) Measured crystal structure of BSCCO with hole doping of $x\approx 22\%$ using STEM adapted from Ref. Song et al. (2019). The red arrows are from the original work, pointing out the oxygen dopants. We show a column of colorful balls for easy identification of the atomic species.
Commencing with undoped Bi-2212, hole doping is incorporated through interstitial oxygen dopants within the material. These additional hole dopants give rise to a diverse range of physical properties including superconductivity, the pseudogap phenomenon, and the presence of shadow bands within the Fermi surface. The subsections below present our findings on various aspects of the hole-doped system at 25% (overdoped): the hole-doped crystal structure including the long-range superlattice modulation and associated placement of the oxygen dopants; the nature of the low-energy magnetic states including spin and charge stripes; and a detailed analysis of the ARPES spectra at the Fermi level including the effects of structural distortions in creating the shadow bands as well as the effect of many-body effects and fluctuating magnetic orders on these spectra.

### v3

# First principles prediction of structural distortions in the cuprates and their impact on electronic structure

@@ Abstract
Materials-realistic microscopic theoretical descriptions of copper-based superconductors are challenging due to their complex crystal structures combined with strong electron interactions. Here, we demonstrate how density functional theory can accurately describe key structural, electronic, and magnetic properties of the normal state of the prototypical cuprate Bi2Sr2CaCu2O8+x (Bi-2212). We emphasize the importance of accounting for energy-lowering structural distortions, which then allows us to: (a) accurately describe the insulating antiferromagnetic (AFM) ground state of the undoped parent compound (in contrast to the metallic state predicted by previous ab initio studies); (b) identify numerous low-energy competing spin and charge stripe orders in the hole-overdoped material nearly degenerate in energy with the AFM ordered state, indicating strong spin fluctuations; (c) predict the lowest-energy hole-doped crystal structure including its long-range structural distortions and oxygen dopant positions that match high-resolution scanning transmission electron microscopy (STEM) measurements; and (d) describe electronic bands near the Fermi energy with flat antinodal dispersions and Fermi surfaces that in agreement with angle-resolved photoemission spectroscopy (ARPES) measurements and provide a clear explanation for the structural origins of the so-called “shadow bands”. We also show how one must go beyond band theory and include fully dynamic spin fluctuations via a many-body approach when aiming to make quantitative predictions to measure the ARPES spectra in the overdoped material. Finally, regarding spatial inhomogeneity, we show that the local structure at the CuO2 layer, rather than dopant electrostatic effects, modulates the local charge-transfer gaps, local correlation strengths, and by extension the local superconducting gaps.
@@ I Introduction
The cuprate superconductors continue to be a fascinating and actively researched class of materials. In addition to their superconducting phases, their normal state has attracted broad research interest due to a wide range of unusual properties. Understanding the physical origin of the AFM insulating phase [1, 2], the pseudogap [3, 4, 5, 6], the flat antinodal dispersion [7, 8, 9, 10], the strange metallicity [11, 12, 13, 14, 15], and the presence of quantum critical fluctuations [16, 17, 18, 19] in the normal state can provide important insights into the underlying mechanisms that can give rise to the superconductivity. Despite plenty of proposed possible mechanisms, including competing orders [20, 21, 22] and preformed pairs [23, 24, 25, 26], the physical origin of this complicated normal state is still unclear.
@@
A comprehensive description of the normal state is challenging due to the strong electronic interactions combined with the structural complexity of typical doped cuprates. While the effect of strong electronic interactions has been extensively studied using accurate many-body methods such as density matrix renormalization group (DMRG) [27, 28] or quantum Monte Carlo (QMC) [29, 30], these studies are often based on idealized effective model Hamiltonian, where the hopping strengths are averaged and symmetrized for simplicity [31, 32, 33]. However, realistic structural distortions in cuprates can result in significant changes to the materials such as additional symmetry breaking [34, 35]. In particular, the structural distortion in cuprates can greatly modify the superconducting gap [36, 37] and local pairing interactions [38, 39, 40, 41, 42, 43, 44, 45]. These properties are missing in the symmetrized effective Hamiltonians. The inclusions of complex lattice distortions of the native material or as introduced by dopants and impurities necessitate a realistic and detailed understanding of the materials from first principles.
@@
Density functional theory [46, 47] (DFT) offers a potent foundational method for investigating the ground-state properties of materials from first principles. For cuprates, DFT has played a pivotal role in constructing effective model Hamiltonians [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. These works studied the high-symmetry cuprate crystal structure using DFT and extracted the low-energy effective model for further quantum many-body studies. However, to ensure the accuracy and realism of these models, it is essential that the DFT calculations capture correctly both the structural properties and the predominant electronic properties of the ground state. Recent DFT studies on transition metal oxides have highlighted the significance of allowing energy-lowering structural distortions to achieve high-quality predictions of materials properties, such as band gaps at eV scale [61, 62, 63].
@@
Bismuth strontium calcium copper oxide Bi2Sr2CaCu2O8+x (BSCCO or Bi-2212) [64] is one of the most intensively studied cuprates and is the focus of in this work. However, prior DFT studies faced challenges due to its intricate structural distortions and superlattice modulations [65, 66, 67, 68]. We demonstrate that by employing modern DFT exchange-correlation functionals and providing an accurate description of the crystalline structure including energy-lowering lattice distortions, we can directly describe the antiferromagnetic insulating ground state of undoped Bi-2212 ($x=0$), the experimentally-observed crystal structure of hole-doped Bi-2212 ($x\approx 0.25$), the presence of competing magnetic and charge stripe orders, as well as crucial details of photoemission spectra such as “shadow bands”. This means that based on DFT calculations, we can correctly ascribe certain experimental observations to specific structural motifs (e.g., shadow bands) while simultaneously helping build microscopically well-justified model Hamiltonians that allow us to study the effects of strong electron correlations further. Our many-body calculations based on such models predict spatial modulations of charge-transfer gaps and correlation strength, consistent with experimental observations [69, 36, 37]. Interestingly, structural modulation of the CuO2 plane is the dominant cause of this phenomenon, far more important than the electrostatic effects from dopant atoms that were assumed to be important in earlier model calculations [38, 39, 41, 42, 45].
@@ II Undoped system
The undoped Bi-2212 possesses a bilayer crystal structure as depicted in Fig. 1(a)(b), where each bilayer comprises two CuO layers separated by one Ca layer and sandwiched between SrO and BiO layers. The smallest unit cell contains 30 atoms and crystallizes in the tetragonal I4/mmm space group, with only one Cu atom in each CuO layer [70, 71, 72]. However, this small unit cell leads to a false non-magnetic ground state due to the artificial assumption of translational invariance of the Cu local moments [73]. In addition, the tetragonal space group of this small unit cell is inconsistent with the orthorhombicity observed in experiments [74]. Here, we start with a larger supercell (60 atoms/cell) with two Cu atoms in each CuO layer, which allows for structural orthorhombicity and the spontaneous symmetry breaking of the magnetic local moments. We ignore the superlattice modulation distortion for now and focus on the distortions in BiO layers. As we will show in the following, the distortions in BiO layers are crucial to raising the antibonding Bi-O band to higher energy.
For simplicity, we first compare different distortion patterns in bulk calculations. After finding the most energetically favorable distortions, we then perform slab calculation to compare with available undoped Bi-2212 experiments such as STM and thin film transports. Because the inter-bilayer van der Waals interactions in Bi-2212 are weak, the slab calculations show very small differences with the bulk results. This is also the reason why many prior studies only focused on slab calculations [75, 76, 77].
@@
To improve the theoretical description, allowing crystal structural distortions in the BiO layers can further lower the total energy of the calculation. Using the conjugate gradient algorithm for structural relaxation, we optimize the structures by initially lowering the symmetry in the BiO layers manually. Fig. 2(a)(b) depicts two typical stable or meta-stable BiO layer patterns whose ground states are both G-AFM ordered on the CuO2 planes. These two distortions have been studied separately in prior studies [80, 81, 82]. Here we will compare the effects of different distortion patterns on electronic structures and demonstrate how the distortions help raise the BiO antibonding band.
@@
Fig. 2(b) shows an orthorhombic pattern, characterized by the lowest symmetry among all three structures presented in Fig. 1 and 2. This pattern further breaks the mirror symmetry of the crystal, resulting in an orthorhombic lattice consistent with observations in experiments [74]. This orthorhombic pattern exhibits the lowest AFM ground state energy, 0.5 eV/Cu lower than the high-symmetry case. It also features the largest local Cu magnetic moments of ±$\pm 0.53\mu_{B}$ as well as the largest BiO bonding gap size of about 1.9 eV among all three Bi-O structural motifs mentioned above.
@@
Structure  hopping (eV)  bond length (Å)  bond angle ()
1 1.0 2.65 90
2 1.8 2.28 112
3 2.3 2.17 93
4 0.9 2.65 93
Table 1: Bi-O hopping strengths, Bi-O bond lengths, and O-Bi-O bond angles in the BiO plane bond for different distorted structures: structure 1 is the high-symmetry structure of Fig. 1(b); 2 is the zigzag distorted structure of Fig. 2(a); 3 is the orthorhombic distorted structure of Fig. 2(b); 4 has the same geometry as structure 3 but with stretched in-plane lattice constant to match the Bi-O bond length of structure 1.
This microscopic picture is consistent with a structural analysis. Generally, hopping strengths depend directly on local structural properties such as bond lengths and angles. We compare four different structures in Table 1: 1. the high-symmetry structure from Fig. 1; 2. the zigzag distorted structure from Fig. 2(a); 3. the orthorhombic distorted structure from Fig. 2(b); 4. the orthorhombic distorted structure with expanded in-plane lattice constant. Structures 1-3 are stable/meta-stable relaxed structures predicted by DFT, and structure 4 is prepared to have the same distortion pattern as structure 3 but the same bond length as the high-symmetry structure 1. To extract the hopping strength, we compute the tight-binding Kohn-Sham Hamiltonian on the maximally localized Wannier basis [83] extracted from our DFT calculations using Wannier90 [84]. The projected Wannier orbitals encompass the $p$-orbitals of Bi and O, as well as the $d$-orbitals of Cu, which are sufficient to describe the bands near the Fermi level [85]. Table 1 lists the Bi-O hopping (tunneling) matrix element for nearest-neighbor in-plane Bi-O pairs for different structures. Upon comparing these four structures, we observe that the bond length has the most significant impact on the hopping strength (shorter bonds give larger hoppings, as expected). The bond angles do influence the hopping strengths but only modestly (compare structures 1 and 4). Notably, the hopping strength in structure 3 (the lowest energy structure predicted by DFT) is the only one large enough to open a band gap by raising the Bi electron pockets (originating from the anti-bonding BiO bands) above the Fermi level, and this leads to the insulating ground state.
@@
We conclude this section by comparing our results with experiments and prior theories. Experimentally, bulk Bi-2212 crystals are conducting due to hole hoping from inevitable excess oxygen $(x>0)$, so comparing our undoped calculations to experiments on the bulk is not fruitful. However, undoped thin films of Bi-2212 ($x=0$) have been realized experimentally: scanning tunneling microscopy (STM) [86, 87, 88] and transport [89, 90] experiments have found an AFM insulating ground state in undoped thin films Bi-2212. (About $x\approx 6\%$ hole-doping is required to turn the insulating Bi-2212 film into a superconductor [89, 90].) Theoretically, some aspects of the undoped ground state have not been described well by prior DFT work which may have hampered further theoretical studies of the effects of doping and other perturbations on the system at a microscopic level. For example, prior DFT works were unable to simultaneously reproduce the AFM order and the insulating behavior of undoped Bi-2212 whether modeling the bulk [73, 81, 82], thin film or or slabs [75, 76, 77], or when using larger more realistic unit cells that include the observed crystal modulation [68, 91]. The prior work found undoped Bi-2212 to be metallic and “self-doped” due to Bi electron pockets at the Fermi surface (pockets that have not been seen in experiments). Some prior works [81, 76, 77] resorted to additional “manual” hole doping to eliminate the finite Bi density of states: while this is a sensible workaround, it would be preferable to have a calculation of the undoped material reflect the insulating experimental system.
@@ III Oxygen-doped system
Figure 4: (a) DFT predicted crystal structure of Bi2Sr2CaCu2O8.25, where the hole doping level is $x=0.25$. Sr, Ca, Cu, Bi, O, and dopant O atoms are marked by green, gray, blue, purple, red, and black balls, respectively. Red arrows further highlight the O dopants. The black solid square in the crystal structure marks the 244-atom unit cell of the doped crystal. The black dashed rectangle highlights the CuO layers in one of the bilayers in the unit cell. (b) Measured crystal structure of BSCCO with hole doping of $x\approx 22\%$ using STEM adapted from Ref. [92]. The red arrows are from the original work, pointing out the oxygen dopants. We show a column of colorful balls for easy identification of the atomic species.
Commencing with undoped Bi-2212, hole doping is incorporated through interstitial oxygen dopants within the material. These additional hole dopants give rise to a diverse range of physical properties including superconductivity, the pseudogap phenomenon, and the presence of Fermi surface. The subsections below present our findings on various aspects of the hole-doped system at 25% (overdoped): the hole-doped crystal structure including the long-range superlattice modulation and associated placement of the oxygen dopants; the nature of the low-energy magnetic states including spin and charge stripes; and a detailed analysis of the ARPES spectra at the Fermi level including the effects of structural distortions in creating the shadow bands as well as the effect of many-body effects and fluctuating magnetic orders on these spectra.